# Universal activated aging and weak ergodicity breaking in spin and structural glasses

### v4

@@ Abstract
I Introduction
@@
Figure 1: Basic aging phenomenon in glasses. Data are obtained for the WCA model at $T=0.25$ with $N=8000$ particles, quenched from $T_{\rm eq}=5$ to $T$. (a) $dE/dt_{\rm w}$ as a function of $t_{\rm w}$ at a few different $T$. The solid line indicates $dE/dt_{\rm w}\sim t_{\rm w}^{-1}$, i.e., $E(t_{\rm w})\sim-\ln(t_{\rm w})$. (b) Non-equilibrium decorrelation time $\tau_{\rm neq}$ as a function of $t_{\rm w}$. The solid line represents $\tau_{\rm neq}\sim t_{\rm w}$. (c) $F(t_{\rm w},t_{\rm w}+t)$ for $t_{\rm w}=0.2,2,20,2\times 10^{2},2\times 10^{3},2\times 10^{4},8\times 10^{5}$ (black curves; from left to right), and $F_{w}(t_{\rm w})$ with $w=1/2$ and $w=3/2$. The dashed line indicates the nonergodicity parameter $f$. (c) $F(t_{\rm w},t_{\rm w}+t)$ plotted as a function of $w=t/t_{\rm w}$ for intermediate waiting times $t_{\rm w}=2,20,2\times 10^{2},2\times 10^{3}$. The blue line is $H$ given by Eq. (2) at the given $\hat{x}=T/T_{\rm WEB}$, and the red line is $\tilde{H}=fH$.
@@
To understand aging in real glassy systems, it is inevitable to go beyond mean-field approximations. The landscape-based approach, which we follow in the current study, is particularly attractive due to its simplicity and deep connections to general stochastic processes modeled by continuous-time random walk  [30, 31]. The trap model is initially proposed by Bouchaud (BTM) [22, 23] to describe aging dynamics in the simplest spin glass model [32], i.e., the random energy model (REM) [33]. An inspiring outcome from the BTM is the so-called weak ergodicity breaking (WEB) in the spin glass phase, where the non-equilibrium correlation function converges, in the large-time limit, to a plateau whose value is determined by the so-called arcsin law [23]. Such asymptotic solutions of the BTM are supported by recent rigorous mathematical derivations [34, 35, 36, 37]. However, this result is inconsistent with Monte Carlo (MC) simulations of the REM (with Gaussian random energies) consisting of a finite number NN of spins  [38, 39]: the obtained aging functions do not show any sign of convergence to the predicted plateau. Obviously, before any rigorous extension can be made, a fundamental task is to reconcile the inconsistency between the theory and MC simulations in the REM. Once this mission is accomplished, the next natural question is about the relevance of trap model predictions to the aging behavior in more realistic glasses [40]. Both problems are addressed in the present study.
@@
II Basic aging phenomenon and weak ergodicity breaking
@@
The aging dynamics can be monitored by one-time observables, such as the potential energy $E(t_{\rm w})$, which typically shows slow decay close to a logarithmic time-dependent function, $E(t_{\rm w})\sim-\ln(t_{\rm w})$ (see Fig. 1a; to be explicit, in Fig. 1 we focus on the simulation results of the WCA model). More informative descriptors are two-time correlation functions, $C(t_{\rm w},t_{\rm w}+t)=\langle A^{*}(t_{\rm w})A(t_{\rm w}+t)\rangle$, where AA is a physical observable, $t_{\rm w}$ the waiting time elapsed after quenching, and $\langle\ldots\rangle$ the average over initial configurations. In structural glasses, one typically takes $A(t)=e^{i{\bf q}\cdot{\bf r}_{i}(t)}$, where ${\bf r}_{i}(t)$ is the position of particle $i$ at time $t$, and $\bf q$} is a wave-vector - then $F_{q}(t_{\rm w},t_{\rm w}+t)=\frac{1}{N}\langle\sum_{i=1}^{N}e^{i{\bf q}\cdot[{\bf r}_{i}(t_{\rm w}+t)-{\bf r}_{i}(t_{\rm w})]}\rangle$ is essentially the non-equilibrium generalization of the standard incoherent scattering function. In spin glasses, it is natural to consider the spin-spin correlations, $C(t_{\rm w},t_{\rm w}+t)=\frac{1}{N}\langle\sum_{i=1}^{N}S_{i}(t_{\rm w})S_{i}(t_{\rm w}+t)\rangle$. Here NN represents the system size, i.e., number of particles or spins. Figure 1b shows that $F_{q}(t_{\rm w},t_{\rm w}+t)$ in the WCA model has a strong dependence on $t_{\rm w}$, which is direct evidence of non-equilibrium dynamics. Following the convention, in this study we focus on a fixed $q=q_{\rm max}$, where $q_{\rm max}$ is the location of the maximum in the structural factor $S(q)$. The subscript of $q$ in $F_{q}$ will be omitted from now on.
Figure 2: Barrier energy distributions in the G-REM and E-REM. (a) A sub-tree formed by local minima and saddle points for the $N=6$ G-REM obtained by the BT algorithm, with corresponding spin configurations and energies ($E$-axis) indicated. (inset) Illustration of two traps: the blue nodes are non-minimum-non-saddle configurations; the barrier energy $\Delta E$ is the energy difference between the local minimum and saddle point. (b) Distributions $p_{\rm G-REM}^{\rm BT}(\Delta E)$ and $p_{\rm E-REM}^{\rm BT}(\Delta E)$ of the G-REM and E-REM by the BT algorithm ($N=20$), compared to $p_{\rm BTM}(\Delta E)$ (Eq. 1) and $p_{\rm GTM}(\Delta E)$ (Eq. 3 with $a=1.12$ and $b=0.96$). (c) Remainder $\ln p(\Delta E)+\Delta E/T_{\rm c}$ for the G-REM (black) and E-REM (red). Open and filled points are obtained by the BT algorithm and TE theory respectively. Lines represent fitting to a Gaussian function. The variance $\sigma^{2}$ and mean $\bar{E}$ obtained by the Gaussian fitting are plotted in (d) as functions of NN. Linear fitting of the data in (d) gives $a=1.12$ and $b=0.96$ ($\sigma^{2}=aN$ and $\bar{E}=bN$).
@@
Figure 3: Aging function $\Pi_{w}(t_{\rm w})$ for the (a,b) E-REM and (c-d) G-REM. In both models, $w=0.5$. Open points are MC simulation data. (a) E-REM results with $N=20$ (points). Data in the plateau regime are fitted by $\Pi(t_{\rm w})=H\left(1+A\,t_{\rm w}^{-\alpha}\right)$ (lines). (b) E-REM data at $T=0.75$. The purple dashed line represents the short-time behavior $\Pi_{w}^{\rm s}(t_{\rm w})=1-Ct_{\rm w}$ with $C\approx 0.214$. The dotted horizontal line represents the BTM plateau $H(w,x)\approx 0.303$. The thermalization time $\tau_{\rm th}=\exp(N/T)$ is marked by crosses. (c) G-REM results with $N=128$ (points). Data at large times are fitted by $\Pi(t_{\rm w})=H\left(1+A\,t_{\rm w}^{-\alpha}\right)-k\ln t_{\rm w}$ (lines). (d) At $T=0.75$, the G-REM data are dominated by logarithmic decay $\Pi(t_{\rm w})\sim-\frac{B}{N}\ln t_{\rm w}$ (lines) in small systems, where $B=0.7$ is a fitting parameter. The horizontal dashed lines represent the GTM plateau $H(w,\hat{x})\approx 0.854$ and the BTM plateau $H(w,x)\approx 0.136$. The fitting parameters $H$ and $\alpha$ for both models are plotted in Fig. 7(a,b).
III Bouchaud’s trap model
@@
Figure 4: WEB in (a,b) E-REM and (c,d) G-REM. In (a,c), $\Pi_{w}(t_{\rm w})$ is obtained with $N=20$ and $w=0.5$ by MC simulations at the rescaled time $\tilde{t}_{\rm w}=Ct_{\rm w}$ with $C=\frac{w\hat{x}}{1+\hat{x}}$. The solid line is $\Pi^{0}_{w}(\tilde{t}_{\rm w})=1-\tilde{t}_{\rm w}$ (see Eq. 5). The dotted horizontal line is $\Pi_{\rm thr}=0.2$. The time scales $\tilde{\tau}^{0}$ (circle) and $\tilde{\tau}$ (cross) are marked. (b,d) $R=\tilde{\tau}/\tilde{\tau}^{0}$ as a function of $T$ for different NN. (inset) $R$ as a function of NN at $T=1.17\,T_{\rm c}$.
@@
In the REM, it was claimed that $T_{\rm c}$ corresponds to the spin glass transition temperature (see Appendix A for the details of the REM). The exponential form Eq. (1) is consistent with mean-field thermodynamic calculations for the glass models in the so-called the one-step replica symmetry breaking (1-RSB) universality class [1]. Interestingly, Eq. (1) is formally identical to the Boltzmann distribution, $P(E)\sim\exp(-E/T)$, of the microstate energy in an equilibrium system - the only difference is the replacement of $T$ by a constant $T_{\rm c}$. Because $P(\Delta E)$ is a static property of the energy landscape, it should be $T$-independent; then a natural guess of the parameter $T_{\rm c}$ in Eq. (1) is the glass transition temperature, which is the only characteristic temperature in simple models such as the REM.
According to Eq. (1) and the Arrhenius law, $\tau$ follows a power-law distribution, $\psi(\tau)\sim\tau^{-(1+x)}$, where $x=T/T_{\rm c}$ is the reduced temperature. Because $0<x<1$ below $T_{\rm c}$, the average trapping time $\langle\tau\rangle$ diverges in the glass phase. Thus the system takes infinite time to reach equilibrium, leading to long-time aging and WEB. Mathematically, WEB is described by the arcsin law [47] (see Appendix B):
@@
Figure 5: Aging in (a-c) WCA and (d-f) amorphous silica models. (a) $F_{w}(t_{\rm w})$ obtained by MD simulations of the WCA model, with $N=16000$. The data in the intermediate time regime are fitted to $F_{w}(t_{\rm w})=\tilde{H}\left(1+A\,t_{\rm w}^{-\alpha}\right)-k\ln t_{\rm w}$ (solid lines). At the two lowest $T$, $\tilde{H}$ is estimated by fitting the plateau (dashed lines), and $\alpha$ cannot be determined due to the presence of a dip. (b) $F_{w}(t_{\rm w})$ at $T=0.26$ with different NN. Inset is the close-up for the data in $1<t_{\rm w}<10^{3}$, fitted to $F_{w}(t_{\rm w})\sim-k\ln t_{\rm w}$ (lines). The parameter $k$ obtained in this way is plotted in (c) as a function of NN at different $T$. The crossover size $n$ is determined by the intersect of two linear fitting lines (when $T<0.35$). (d-e) Similar plots for amorphous silica (${\rm SiO}_{2}$). In (d,f), $N=31944$; in (e), $T=4000$ K. In all panels, $w=1/2$. The parameters $H$ and $\alpha$ are plotted in Fig. 7(a,b) for both models.
IV Generalized trap model
To understand why the BTM fails to describe the simulation results, one should naturally re-examine the underlying assumptions. Based on the numerical data obtained for REMs, we have examined three major assumptions as described above: (i) the Arrhenius law $\tau\sim\exp(\Delta E/T)$, (ii) the renewal mechanism $\psi(\tau_{1},\tau_{2})=\psi(\tau_{1})\psi(\tau_{2})$, and (iii) the exponential barrier distribution Eq. (1). We find that the first two assumptions hold (see Appendix A), but the third one requires correction (see below). Based on these observations, we propose a GTM, which retains all assumptions of the original BTM, except for the exponential barrier distribution. In the GTM, we add a Gaussian correction term to the barrier distribution,
@@ IV Generalized trap model
where $a$ is a model-dependent parameter. The energy scale $\bar{E}=bN$ is assumed to be extensive, where $b$ is also model-dependent. Note that the Gaussian term, which follows the standard central limit theorem, is essentially caused by finite-size effects. Equation (3) will be explicitly examined in the REMs below, but for now, we take it as input and derive the corresponding aging function theoretically (see details in the Supplemental Material, SM, Sec. S1).
@@
Expanding Eq. (3) gives (neglecting the constant term), $\ln p_{\rm GTM}(\Delta E)\sim-\frac{1}{T_{\rm c}}\left(1-\frac{\bar{E}T_{c}}{Na}\right)\Delta E-\frac{1}{2Na}\Delta E^{2}$. The linear coefficient suggests an effective $x$ parameter, $\hat{x}=\left(1-bT_{\rm c}/a\right)x=(1-W)T/T_{\rm c}$, where the parameter $W=bT_{\rm c}/a$ weights the contributions of linear and Gaussian terms in Eq. (3). Importantly, $W$ is independent of NN, and thus this modification does not disappear even in the thermodynamic limit. However, if we take $N\to\infty$ first in Eq. (3), then obviously the GTM degenerates with the BTM. It suggests that the double limits $N\to\infty$ and $t_{\rm w}\to\infty$ are not interchangeable, once WEB occurs. In the rigorous analyses performed previously [34, 35, 36, 37], the order is $N\to\infty$ first and then $t_{\rm w}\to\infty$. However, in MC simulations, generally long-time simulations ($t_{\rm w}\to\infty$) are performed for small NN, then finite-size analysis is performed to extrapolate the behavior in the thermodynamic limit. In the latter case, the correction due to the Gaussian term is non-negligible even in large systems, which explains the discrepancy between simulation and rigorous results.
@@
The second term $\frac{B}{N}\ln t_{\rm w}$ implies that, when NN is small, this logarithmic decay will destroy the asymptotic plateau - this is what we observe in simulations of G-REM.
@@
Figure 6: Static lengths. The following data are plotted: $\xi_{\rm ag}$ for the WCA model (this work), $\xi_{\rm Hessian}$ for the KA-LJ model from [43] ($T$ is rescaled by a factor of $3/4$), $\xi_{\rm PTS}$ for the WCA model from [42], and $\xi_{\rm PTS}$ for HSs from  [50]. The HS $\xi_{\rm PTS}$ is plotted as a function of $(Z^{-1}-Z^{-1}_{\rm K})/(Z_{\rm MCT}^{-1}-Z^{-1}_{\rm K})$. The black line represents the fitting $\xi\sim(T-T_{\rm K})^{-\nu}$, where $T_{\rm K}=0.21$ and $\nu=0.61(1).$ (inset) $\xi$ as a function of $T$.
@@
Figure 7: Unified phase diagram of activated aging and WEB in glasses. (a) WEB order parameter $H$ and (b) power-law convergence exponent $\alpha$ obtained by simulations of four glass models ($w=t/t_{\rm w}=1/2$). The theoretical lines are Eqs. (2) and (6). (c) In the unified phase diagram, WEB and ergodic phases are separated by the $T/T_{\rm SEB}=1/(1-W)$ line, and the $H$-axis quantifies the order parameter. Each glass model corresponds to one line in the plot with the corresponding $W$. Insets are schematic energy landscapes in ergodic and WEB phases; the arrow indicates hopping to lower energy basins during aging.
V Aging in spin glass random energy models
@@
In REMs, we examine both the assumptions and predictions of the GTM. We first focus on the assumptions. As shown in Appendix A, data obtained from numerical simulations are consistent with the Arrhenius law and renewal mechanism. Next we examine the barrier energy distribution, Eq. (3), using two independent approaches. The first one is an exhaustive enumeration of energy barriers using the barrier-tree (BT) algorithm [51, 52, 53, 54] (see Fig. 2a for an example of a sub-tree and Appendix A for algorithm details). The BT method provides the exact distribution $p^{\rm BT}(\Delta E)$, but is restricted to small systems ($N\leq 20$). We also develop a tree-expansion (TE) theory to compute the distribution analytically and approximately, giving $p^{\rm TE}(\Delta E)$ for any NN (see SM Sec. S2). The difference between $p^{\rm BT}(\Delta E)$ and $p^{\rm TE}(\Delta E)$ is unnoticeable, as shown in Fig 2(b-d).
@@
While the distribution $p_{\rm E-REM}^{\rm BT}(\Delta E)$ of the E-REM obtained by the BT algorithm is very close to the exponential form $p_{\rm BTM}(\Delta E)$ (see Eq. 1), it is evident that the G-REM distribution $p_{\rm G-REM}^{\rm BT}(\Delta E)$ deviates from a pure exponential distribution (Fig. 2b). To analyze this deviation more carefully, the exponential part Eq. (1) is subtracted from the full distribution, and the remainder $\ln[p^{\rm BT}(\Delta E)]+\Delta E/T_{\rm c}$ is plotted in Fig. 2c for both models. The data for G-REM in Fig. 2c can be well fitted by a quadratic function $f(\Delta E)=c-\frac{(\Delta E-\bar{E})^{2}}{2\sigma^{2}}$. The variance $\sigma^{2}=aN$ and mean $\bar{E}=bN$ linearly depend on NN, as shown in Fig. 2d. From the linear fitting, we obtain $a=1.12$ and $b=0.96$ for the G-REM. In contrast, the remainder of the E-REM in Fig. 2c is negligible. Thus for the E-REM, the barrier distribution is well described by a pure exponential function as in the BTM.

### v5

# thanks: These two authors contributed equallythanks: These two authors contributed equally

@@ Abstract
Teaser: Weak ergodicity breaking occurs prior to strong ergodicity breaking in spin and structural glasses.
Introduction
@@
FIG. 1: Basic aging phenomenon in glasses. Data are obtained for the WCA model at $T=0.25$ with $N=8000$ particles, quenched from $T_{\rm eq}=5$ to $T$. (a) $dE/dt_{\rm w}$ as a function of $t_{\rm w}$ at a few different $T$. The solid line indicates $dE/dt_{\rm w}\sim t_{\rm w}^{-1}$, i.e., $E(t_{\rm w})\sim-\ln(t_{\rm w})$. (b) Non-equilibrium decorrelation time $\tau_{\rm neq}$ as a function of $t_{\rm w}$. The solid line represents $\tau_{\rm neq}\sim t_{\rm w}$. (c) $F(t_{\rm w},t_{\rm w}+t)$ for $t_{\rm w}=0.2,2,20,2\times 10^{2},2\times 10^{3},2\times 10^{4},8\times 10^{5}$ (black curves; from left to right), and $F_{w}(t_{\rm w})$ with $w=1/2$ and $w=3/2$. The dashed line indicates the nonergodicity parameter $f$. (c) $F(t_{\rm w},t_{\rm w}+t)$ plotted as a function of $w=t/t_{\rm w}$ for intermediate waiting times $t_{\rm w}=2,20,2\times 10^{2},2\times 10^{3}$. The blue line is $H$ given by Eq. (2) at the given $\hat{x}=T/T_{\rm WEB}$, and the red line is $\tilde{H}=fH$.
@@
To understand aging in real glassy systems, it is inevitable to go beyond mean-field approximations. The landscape-based approach, which we follow in the current study, is particularly attractive due to its simplicity and deep connections to general stochastic processes modeled by continuous-time random walk  [30, 31]. The trap model is initially proposed by Bouchaud (BTM) [22, 23] to describe aging dynamics in the simplest spin glass model [32], i.e., the random energy model (REM) [33]. An inspiring outcome from the BTM is the so-called weak ergodicity breaking (WEB) in the spin glass phase, where the non-equilibrium correlation function converges, in the large-time limit, to a plateau whose value is determined by the so-called arcsin law [23]. Such asymptotic solutions of the BTM are supported by recent rigorous mathematical derivations [34, 35, 36, 37]. However, this result is inconsistent with Monte Carlo (MC) simulations of the REM (with Gaussian random energies) consisting of a finite number $N$ of spins  [38, 39]: the obtained aging functions do not show any sign of convergence to the predicted plateau. Obviously, before any rigorous extension can be made, a fundamental task is to reconcile the inconsistency between the theory and MC simulations in the REM. Once this mission is accomplished, the next natural question is about the relevance of trap model predictions to the aging behavior in more realistic glasses [40]. Both problems are addressed in the present study.
@@
Results
Basic aging phenomenon and weak ergodicity breaking
@@ Results
The aging dynamics can be monitored by one-time observables, such as the potential energy $E(t_{\rm w})$, which typically shows slow decay close to a logarithmic time-dependent function, $E(t_{\rm w})\sim-\ln(t_{\rm w})$ (see Fig. 1a; to be explicit, in Fig. 1 we focus on the simulation results of the WCA model). More informative descriptors are two-time correlation functions, $C(t_{\rm w},t_{\rm w}+t)=\langle A^{*}(t_{\rm w})A(t_{\rm w}+t)\rangle$, where $A$ is a physical observable, $t_{\rm w}$ the waiting time elapsed after quenching, and $\langle\ldots\rangle$ the average over initial configurations. In structural glasses, one typically takes $A(t)=e^{i{\bf q}\cdot{\bf r}_{i}(t)}$, where ${\bf r}_{i}(t)$ is the position of particle $i$ at time $t$, and $\bf q$} is a wave-vector - then $F_{q}(t_{\rm w},t_{\rm w}+t)=\frac{1}{N}\langle\sum_{i=1}^{N}e^{i{\bf q}\cdot[{\bf r}_{i}(t_{\rm w}+t)-{\bf r}_{i}(t_{\rm w})]}\rangle$ is essentially the non-equilibrium generalization of the standard incoherent scattering function. In spin glasses, it is natural to consider the spin-spin correlations, $C(t_{\rm w},t_{\rm w}+t)=\frac{1}{N}\langle\sum_{i=1}^{N}S_{i}(t_{\rm w})S_{i}(t_{\rm w}+t)\rangle$. Here $N$ represents the system size, i.e., number of particles or spins. Figure 1b shows that $F_{q}(t_{\rm w},t_{\rm w}+t)$ in the WCA model has a strong dependence on $t_{\rm w}$, which is direct evidence of non-equilibrium dynamics. Following the convention, in this study we focus on a fixed $q=q_{\rm max}$, where $q_{\rm max}$ is the location of the maximum in the structural factor $S(q)$. The subscript of $q$ in $F_{q}$ will be omitted from now on.
FIG. 2: Aging function $\Pi_{w}(t_{\rm w})$ for the (a,b) E-REM and (c-d) G-REM. In both models, $w=0.5$. Open points are MC simulation data. (a) E-REM results with $N=20$ (points). Data in the plateau regime are fitted by $\Pi(t_{\rm w})=H\left(1+A\,t_{\rm w}^{-\alpha}\right)$ (lines). (b) E-REM data at $T=0.75$. The purple dashed line represents the short-time behavior $\Pi_{w}^{\rm s}(t_{\rm w})=1-Ct_{\rm w}$ with $C\approx 0.214$. The dotted horizontal line represents the BTM plateau $H(w,x)\approx 0.303$. The thermalization time $\tau_{\rm th}=\exp(N/T)$ is marked by crosses. (c) G-REM results with $N=128$ (points). Data at large times are fitted by $\Pi(t_{\rm w})=H\left(1+A\,t_{\rm w}^{-\alpha}\right)-k\ln t_{\rm w}$ (lines). (d) At $T=0.75$, the G-REM data are dominated by logarithmic decay $\Pi(t_{\rm w})\sim-\frac{B}{N}\ln t_{\rm w}$ (lines) in small systems, where $B=0.7$ is a fitting parameter. The horizontal dashed lines represent the GTM plateau $H(w,\hat{x})\approx 0.854$ and the BTM plateau $H(w,x)\approx 0.136$. The fitting parameters $H$ and $\alpha$ for both models are plotted in Fig. 7(a,b).
@@
FIG. 3: Barrier energy distributions in the G-REM and E-REM. (a) A sub-tree formed by local minima and saddle points for the $N=6$ G-REM obtained by the BT algorithm, with corresponding spin configurations and energies ($E$-axis) indicated. (inset) Illustration of two traps: the blue nodes are non-minimum-non-saddle configurations; the barrier energy $\Delta E$ is the energy difference between the local minimum and saddle point. (b) Distributions $p_{\rm G-REM}^{\rm BT}(\Delta E)$ and $p_{\rm E-REM}^{\rm BT}(\Delta E)$ of the G-REM and E-REM by the BT algorithm ($N=20$), compared to $p_{\rm BTM}(\Delta E)$ (Eq. 1) and $p_{\rm GTM}(\Delta E)$ (Eq. 3 with $a=1.12$ and $b=0.96$). (c) Remainder $\ln p(\Delta E)+\Delta E/T_{\rm c}$ for the G-REM (black) and E-REM (red). Open and filled points are obtained by the BT algorithm and TE theory respectively. Lines represent fitting to a Gaussian function. The variance $\sigma^{2}$ and mean $\bar{E}$ obtained by the Gaussian fitting are plotted in (d) as functions of $N$. Linear fitting of the data in (d) gives $a=1.12$ and $b=0.96$ ($\sigma^{2}=aN$ and $\bar{E}=bN$).
Bouchaud’s trap model
@@
FIG. 4: WEB in (a,b) E-REM and (c,d) G-REM. In (a,c), $\Pi_{w}(t_{\rm w})$ is obtained with $N=20$ and $w=0.5$ by MC simulations at the rescaled time $\tilde{t}_{\rm w}=Ct_{\rm w}$ with $C=\frac{w\hat{x}}{1+\hat{x}}$. The solid line is $\Pi^{0}_{w}(\tilde{t}_{\rm w})=1-\tilde{t}_{\rm w}$ (see Eq. 5). The dotted horizontal line is $\Pi_{\rm thr}=0.2$. The time scales $\tilde{\tau}^{0}$ (circle) and $\tilde{\tau}$ (cross) are marked. (b,d) $R=\tilde{\tau}/\tilde{\tau}^{0}$ as a function of $T$ for different NN. (inset) $R$ as a function of NN at $T=1.17\,T_{\rm c}$.
@@
In the REM, it was claimed that $T_{\rm c}$ corresponds to the spin glass transition temperature (see Materials and Methods A for the details of the REM). The exponential form Eq. (1) is consistent with mean-field thermodynamic calculations for the glass models in the so-called the one-step replica symmetry breaking (1-RSB) universality class [1]. Interestingly, Eq. (1) is formally identical to the Boltzmann distribution, $P(E)\sim\exp(-E/T)$, of the microstate energy in an equilibrium system - the only difference is the replacement of $T$ by a constant $T_{\rm c}$. Because $P(\Delta E)$ is a static property of the energy landscape, it should be $T$-independent; then a natural guess of the parameter $T_{\rm c}$ in Eq. (1) is the glass transition temperature, which is the only characteristic temperature in simple models such as the REM.
According to Eq. (1) and the Arrhenius law, $\tau$ follows a power-law distribution, $\psi(\tau)\sim\tau^{-(1+x)}$, where $x=T/T_{\rm c}$ is the reduced temperature. Because $0<x<1$ below $T_{\rm c}$, the average trapping time $\langle\tau\rangle$ diverges in the glass phase. Thus the system takes infinite time to reach equilibrium, leading to long-time aging and WEB. Mathematically, WEB is described by the arcsin law [47]:
@@
FIG. 5: Aging in (a-c) WCA and (d-f) amorphous silica models. (a) $F_{w}(t_{\rm w})$ obtained by MD simulations of the WCA model, with $N=16000$. The data in the intermediate time regime are fitted to $F_{w}(t_{\rm w})=\tilde{H}\left(1+A\,t_{\rm w}^{-\alpha}\right)-k\ln t_{\rm w}$ (solid lines). At the two lowest $T$, $\tilde{H}$ is estimated by fitting the plateau (dashed lines), and $\alpha$ cannot be determined due to the presence of a dip. (b) $F_{w}(t_{\rm w})$ at $T=0.26$ with different NN. Inset is the close-up for the data in $1<t_{\rm w}<10^{3}$, fitted to $F_{w}(t_{\rm w})\sim-k\ln t_{\rm w}$ (lines). The parameter $k$ obtained in this way is plotted in (c) as a function of NN at different $T$. The crossover size $n$ is determined by the intersect of two linear fitting lines (when $T<0.35$). (d-e) Similar plots for amorphous silica (${\rm SiO}_{2}$). In (d,f), $N=31944$; in (e), $T=4000$ K. In all panels, $w=1/2$. The parameters $H$ and $\alpha$ are plotted in Fig. 7(a,b) for both models.
Generalized trap model
To understand why the BTM fails to describe the simulation results, one should naturally re-examine the underlying assumptions. Based on the numerical data obtained for REMs, we have examined three major assumptions as described above: (i) the Arrhenius law $\tau\sim\exp(\Delta E/T)$, (ii) the renewal mechanism $\psi(\tau_{1},\tau_{2})=\psi(\tau_{1})\psi(\tau_{2})$, and (iii) the exponential barrier distribution Eq. (1). We find that the first two assumptions hold, but the third one requires correction (see below). Based on these observations, we propose a GTM, which retains all assumptions of the original BTM, except for the exponential barrier distribution. In the GTM, we add a Gaussian correction term to the barrier distribution,
@@
where $a$ is a model-dependent parameter. The energy scale $\bar{E}=bN$ is assumed to be extensive, where $b$ is also model-dependent. Note that the Gaussian term, which follows the standard central limit theorem, is essentially caused by finite-size effects. Equation (3) will be explicitly examined in the REMs below, but for now, we take it as input and derive the corresponding aging function theoretically (see details in the Supplementary Materials, SM, Sec. S1).
@@
Expanding Eq. (3) gives (neglecting the constant term), $\ln p_{\rm GTM}(\Delta E)\sim-\frac{1}{T_{\rm c}}\left(1-\frac{\bar{E}T_{c}}{Na}\right)\Delta E-\frac{1}{2Na}\Delta E^{2}$. The linear coefficient suggests an effective $x$ parameter, $\hat{x}=\left(1-bT_{\rm c}/a\right)x=(1-W)T/T_{\rm c}$, where the parameter $W=bT_{\rm c}/a$ weights the contributions of linear and Gaussian terms in Eq. (3). Importantly, $W$ is independent of $N$, and thus this modification does not disappear even in the thermodynamic limit. However, if we take $N\to\infty$ first in Eq. (3), then obviously the GTM degenerates with the BTM. It suggests that the double limits $N\to\infty$ and $t_{\rm w}\to\infty$ are not interchangeable, once WEB occurs. In the rigorous analyses performed previously [34, 35, 36, 37], the order is $N\to\infty$ first and then $t_{\rm w}\to\infty$. However, in MC simulations, generally long-time simulations ($t_{\rm w}\to\infty$) are performed for small $N$, then finite-size analysis is performed to extrapolate the behavior in the thermodynamic limit. In the latter case, the correction due to the Gaussian term is non-negligible even in large systems, which explains the discrepancy between simulation and rigorous results.
@@
The second term $\frac{B}{N}\ln t_{\rm w}$ implies that, when $N$ is small, this logarithmic decay will destroy the asymptotic plateau - this is what we observe in simulations of G-REM.
@@
FIG. 6: Static lengths. The following data are plotted: $\xi_{\rm ag}$ for the WCA model (this work), $\xi_{\rm Hessian}$ for the KA-LJ model from [43] ($T$ is rescaled by a factor of $3/4$), $\xi_{\rm PTS}$ for the WCA model from [42], and $\xi_{\rm PTS}$ for HSs from  [50]. The HS $\xi_{\rm PTS}$ is plotted as a function of $(Z^{-1}-Z^{-1}_{\rm K})/(Z_{\rm MCT}^{-1}-Z^{-1}_{\rm K})$. The black line represents the fitting $\xi\sim(T-T_{\rm K})^{-\nu}$, where $T_{\rm K}=0.21$ and $\nu=0.61(1).$ (inset) $\xi$ as a function of $T$.
@@
FIG. 7: Unified phase diagram of activated aging and WEB in glasses. (a) WEB order parameter $H$ and (b) power-law convergence exponent $\alpha$ obtained by simulations of four glass models ($w=t/t_{\rm w}=1/2$). The theoretical lines are Eqs. (2) and (6). (c) In the unified phase diagram, WEB and ergodic phases are separated by the $T/T_{\rm SEB}=1/(1-W)$ line, and the $H$-axis quantifies the order parameter. Each glass model corresponds to one line in the plot with the corresponding $W$. Insets are schematic energy landscapes in ergodic and WEB phases; the arrow indicates hopping to lower energy basins during aging.
Aging in spin glass random energy models
@@
In REMs, we examine both the assumptions and predictions of the GTM. We first focus on the assumptions. As shown in SM Sec. S2, data obtained from numerical simulations are consistent with the Arrhenius law and renewal mechanism. Next we examine the barrier energy distribution, Eq. (3), using two independent approaches. The first one is an exhaustive enumeration of energy barriers using the barrier-tree (BT) algorithm [51, 52, 53, 54] (see Fig. 3a for an example of a sub-tree and Materials and Methods A for algorithm details). The BT method provides the exact distribution $p^{\rm BT}(\Delta E)$, but is restricted to small systems ($N\leq 20$). We also develop a tree-expansion (TE) theory to compute the distribution analytically and approximately, giving $p^{\rm TE}(\Delta E)$ for any $N$ (see SM Sec. S3). The difference between $p^{\rm BT}(\Delta E)$ and $p^{\rm TE}(\Delta E)$ is unnoticeable, as shown in Fig 3(b-d).
@@
While the distribution $p_{\rm E-REM}^{\rm BT}(\Delta E)$ of the E-REM obtained by the BT algorithm is very close to the exponential form $p_{\rm BTM}(\Delta E)$ (see Eq. 1), it is evident that the G-REM distribution $p_{\rm G-REM}^{\rm BT}(\Delta E)$ deviates from a pure exponential distribution (Fig. 3b). To analyze this deviation more carefully, the exponential part Eq. (1) is subtracted from the full distribution, and the remainder $\ln[p^{\rm BT}(\Delta E)]+\Delta E/T_{\rm c}$ is plotted in Fig. 3c for both models. The data for G-REM in Fig. 3c can be well fitted by a quadratic function $f(\Delta E)=c-\frac{(\Delta E-\bar{E})^{2}}{2\sigma^{2}}$. The variance $\sigma^{2}=aN$ and mean $\bar{E}=bN$ linearly depend on $N$, as shown in Fig. 3d. From the linear fitting, we obtain $a=1.12$ and $b=0.96$ for the G-REM. In contrast, the remainder of the E-REM in Fig. 3c is negligible. Thus for the E-REM, the barrier distribution is well described by a pure exponential function as in the BTM.